# Foundation Models for Music: A Survey

### v1

@@ Abstract
In recent years, foundation models (FMs) such as large language models (LLMs) and latent diffusion models (LDMs) have profoundly impacted diverse sectors, including music. This comprehensive review examines state-of-the-art (SOTA) pre-trained models and foundation models in music, spanning from representation learning, generative learning and multimodal learning. We first contextualise the significance of music in various industries and trace the evolution of AI in music. By delineating the modalities targeted by foundation models, we discover many of the music representations are underexplored in FM development. Then, emphasis is placed on the lack of versatility of previous methods on diverse music applications, along with the potential of FMs in music understanding, generation and medical application. By comprehensively exploring the details of the model pre-training paradigm, architectural choices, tokenisation, finetuning methodologies and controllability, we emphasise the important topics that should have been well explored, like instruction tuning and in-context learning, scaling law and emergent ability, as well as long-sequence modelling etc. A dedicated section presents insights into music agents, accompanied by a thorough analysis of datasets and evaluations essential for pre-training and downstream tasks. Finally, by underscoring the vital importance of ethical considerations, we advocate that following research on FM for music should focus more on such issues as interpretability, transparency, human responsibility, and copyright issues. The paper offers insights into future challenges and trends on FMs for music, aiming to shape the trajectory of human-AI collaboration in the music realm.
@@ III-A1 Traditional Music Information Retrieval Tasks
Given the extensive task list above, there is no single unified resource for providing task definitions and benchmark methodologies, although there exist textbooks which cover a subset of the above tasks (e.g. [Mül15, Ler22]), with ISMIR conference proceedings [^10] https://ismir.net/conferences/ being useful for providing additional technical details. Review and overview papers do exist for specific tasks (which might or might not be up to date with related literature depending on the publication year), including on onset detection [BDA+05], music transcription [BDDE19], chord estimation [MSRNDB14], melody extraction [SGER14], score following [OLS03], music source separation [CFL+18], music classification [FLTZ10], audio identification [CBKH05], cover song detection [SGH10], optical music recognition [CZJP20], plus general roadmaps for the community [SMB+13].
@@ III-A2 Multimodal Music Understanding Tasks
Music instruction following Mu-llama [LHSS23a] and MusiLingo [DML+24] integrated MERT[LYZ+24], a pre-trained music encoder, and Llama, a pre-trained natural language decoder. Mu-llama employs a Llama-adapter on the upper transformer layers for modality alignment [ZHL+23a] while MusiLingo utilises a simple linear projection on the initial transformer layer, yielding enhanced performance. Given the diverse range of tasks within the Music Understanding category, recent works have explored the application of Multimodal Instruction Tuning to enhance the performance of Music Language Models. LLark [GDSB23b] and JMLA [DYL+24] investigate how this approach can be effectively applied to the training of Music LLMs, aiming to improve their versatility and effectiveness across a wide spectrum of music-related tasks. Both models employ a similar architectural approach, utilizing cross-attention modules to connect audio encoders with pre-trained LLMs. This integration allows for effective processing of both audio and textual information. To create robust instruction tuning datasets, these studies leverage traditional music tagging datasets, employing data augmentation techniques to expand them into comprehensive datasets suitable for fine-tuning Multimodal LLMs. By incorporating this approach, these models demonstrate the potential for more adaptive and comprehensive music understanding capabilities, effectively bridging the gap between audio processing and natural language understanding in the music domain.
@@ III-A3 Vocal Music Understanding
There have been several attempts to approach singer ID through representation learning. [LN19] proposed a triplet network that learns a joint embedding for monophonic and mixture singing. [ZQY+21] transformed the classification problem into a query searching task by replacing the softmax layer with a k-nearest neighbors (KNN) layer. [TLR23] trained SSL features on source-separated singing voice and compared them with speech SSL features in terms of classification accuracy.
@@ IV-F Foundation Models as Music Agents
Broadly speaking, an agent is a kind of entities that have desires, beliefs, intentions, and the ability to take actions [ZNAP95]. In the realm of AI, an agent is an artificial entity that can set objectives autonomously or semi-autonomously, perceive the environment, work out plans, reason and make decisions, and take actions to achieve these objectives, powered by the technologies of artificial intelligence [XCG+23]. In terms of music agents, they refer to AI agents dedicated to music and related domains, which can understand, generate, and convert music in symbolic, acoustic, and other modalities.
The development of AI agents has undergone a long process. Early attempts at building AI agents focus on improving the specific capabilities of an agent, such as sensing the environment, symbolic reasoning, taking actions with reinforcement learning, or handling specific tasks. However, the key to AI agents is the general capabilities of reasoning, planning, perception, and action. While focusing on specific aspects of AI agents can make progress steadily, it cannot make fundamental improvements to push AI agents to practical usage.
Large language models (LLMs) that emerged in recent years have revolutionised the era of language and multimodality, and brought strong capabilities in reasoning, planning, perception, and generation, which make them a good fit as the core module of AI agents. Some pioneer investigations such as AutoGPT [Teab], HuggingGPT [SST+23a], and Visual ChatGPT [WYQ+23] leverage LLMs to understand user requests, reason and plan according to objectives, decompose complicated tasks into subtasks, revoke external tools for task execution, and response generation. On the one hand, LLM-based agents extend the capabilities of LLMs to handle complicated and/or multimodality tasks. On the other hand, LLMs empowered AI agents to a new level with strong general capabilities that have never been possible using traditional AI agent technologies. Due to the great potential of LLM-powered AI agents, a lot of research efforts have been made in this area, such as improving the reasoning and planning capabilities [WWS+22, Teaa], benchmarking AI agents [SST+23b], enhancing tool use [SDYD+23, YSC+24], designing agents to handle challenging tasks [SST+23a], and applying AI agents to more domains, including audio [LZL+23, LZL+24] and music [YSL+23]. In the next paragraph, we will review the typical AI agent in music and audio domains.
@@
  • Loop Copilot [ZMX+23] introduces an innovative system combining large language models with specialised AI music models to streamline the collaborative creation of music loops. This system utilises a conversational interface for dynamic, iterative music editing, and a global attribute table to ensure consistency throughout the creation process. Besides, it is not limited to creating music based on vague text inputs but allows for fine-grained musical edits, including adding orremoving tracks, and making localised adjustments to modes and tempos. This capability enhances the system’s utility in detailed music production tasks.
@@
Although there are some preliminary investigations on music agents, there are still a lot of space to improve. We introduce the some future work of music agents as follows.
  • Improve reasoning and planning ability. The key capability of the music agents is enabled by the foundation model behind. When handling complex tasks, the reasoning and planning ability of the foundation model plays a critical role in understanding the user request, decomposing the whole task into subtasks, organizing complicated working flows, and calling suitable models and tool for task execution [HC23, HGM+23]. Thus, improving the reasoning and planning capability has always been the top priorities in building music agents as well as general artificial intelligence.
  • From semi-autonomous to full-autonomous agents. Nowadays, most music agents are semi-autonomous, requiring a humans in the loop, such as initiating user requests, setting system configurations, providing suitable model and tool sets, and strong interactions in the agent workflows. While, Autonomous agents based on LLM are popular research directionss[HGM+23, WMF+24, YLL+24, YLY+24]. In order to play more important and critical roles in helping musician and music consumers, music agents should evolve from semi-autonomous to full-autonomous, gradually reducing the degree of user interference when fulfilling complex user requests.
@@ IV-G Scaling Laws for Music
Scaling laws [KMH+20, HKK+20, HBM+22, AAA+23] characterise the performance of large language models (LLMs) (usually cross-entropy loss on a held-out set) scales as a power law with model size, dataset size, as well as training compute. Scaling laws can predict the loss of LLMs with a given number of model parameters and data size (the number of training tokens, i.e., batch size * training steps) and can determine the optimal configuration of model size and number of training tokens given a computational budget. There are different formulations for scaling laws, such as OpenAI’s scaling laws [KMH+20], Chinchilla scaling laws [HBM+22], and data-constraint scaling laws [MRB+24]. OpenAI’s scaling laws are more precise, while Chinchilla’s laws are simplified based on some assumptions. For simplicity [^14] Strictly speaking, Chinchilla scaling laws have obvious errors. For example, given two models with model size $N_{1}$ and $N_{2}$ where $N_{1}>N_{2}$, when increasing the same amount of training tokens $\Delta D$, the loss reductions of the two models are the same according to Chinchilla laws as shown in Equation 2. However, it is evidently not the case in practice since the larger model $N_{1}$ will have more loss reduction than $N_{2}$. But for ease of explanation, we just use Chinchilla scaling laws as an example here., we briefly describe Chinchilla scaling laws as follows:
@@
There are several typical usage of scaling laws: 1) predict the loss of LLMs with a given model size and training tokens; 2) allocate the optimal model/data size given a compute budget; 3) determine the relative scaling of model and data with an increased compute budget; 3) determine the ratio of data to model given a compute budget.
Traditional scaling laws are mostly studied in language language models for text [KMH+20, HKK+20, HBM+22, AAA+23, MRB+24], which are not applicable for music. MuPT [QBM+24] is first work to study the scaling laws for music, with dedicated design on symbolic music data representations and scaling laws in the data-constraint and over-fitting scenarios. While the great benefits of MuPT in guiding the modelling scaling of symbolic music generation, there is still a lack of scaling laws for music understanding and generation in the audio domain. Later, there are some studies on scaling laws for multimodality [HKK+20, AYC+23] and especially for speech-language model [CM24]. Since multimodal data such as images, videos, and speech are high-dimensional and complex and not usually modelled by language models, it is not straightforward to study the scaling laws on multimodal data. Therefore, they usually tokenise multimodal data into tokens and model them as the next token prediction as in large language models. In this way, they could study the scaling laws similarly to those in text-based large language models.
@@
Since there is not much research on scaling laws for music in the audio domain (although the scaling laws from multimodal and especially speech domain could provide much insight into the audio music domain, it is not directly applicable), and the emergent ability of music FMs is under-explored, we call for actions to proactively study the scaling properties of different data representations and model paradigms for audio music understanding and generation in the future. Such research topics include but are not limited to which type of downstream capability of large models can be inferred by the small models, how the properties of audio tokenisers relate to the parameters of scaling law, and what the minimal requirement of the model’s parameters number or training volumes for specific type of downstream tasks etc.
@@ IV-H Additional Future Improvements
In this subsection, we will shortly discuss potential research topics on foundation models for music. Other than the advanced techniques of foundation models discussed in previous subsections, the following paragraphs will include domain knowledge, long sequence modelling and causal modelling. For the fairness, transparency, and potential issue of model bias, please refer to subsectionVI-A2.
@@ IV-H1 Domain Knowledge
The model architecture described in this section so far, apart from small customisation, can be construed as ‘standard’ in the sense that it has been explored for a plethora of other tasks both somewhat linked with music (e.g. speech) as well as entirely different from it (e.g. machine translation). Although such an approach is quite appealing [HBL12], whether it is limited quantities of data or inadequacies of architecture/optimisation have led to a substantial interest in incorporating domain knowledge. This reversal towards an appreciation of domain expertise has brought about a revival in domain-specific engineering of machine learning approaches in many areas, including music [Ser13]. There are quite a few domains where domain knowledge for music foundation models can be sourced from. The acoustic domain gave rise to spectrum-based representations, which continue to find their use in deep learning architectures [DJP+20a]. The same applies to psychoacoustic chroma tokens [M1̈5] which can assist deep learning should it fail to automatically infer the importance of timbre information from the data [LYZ+24]. Music theory provides a further rich source of information which has inspired numerous approaches, and we believe we will continue to do so.
@@
One key challenge here is how to integrate music theory and related concepts into existing standard architectures rather than designing custom architectures to fit one or more music theories, as was done before [Wid98] and still remains popular in the area [HOB23], or uncover new ones by means of automatic theorists [YV16]. One popular direction for integrating music theory into Transformer and related architectures is input modification. Although a simple approach, it provides opportunities for encoding a wide range of music-theoretical information. So far, the work in this direction has focused on encoding lower-level relative attributes, such as pitch interval, duration and onset [GKH23], and some higher-level attributes, such as bars [WERA16]. There remains to be an extension of this work to other higher-level relative attributes. The same applies to absolute attributes such as melody, harmony and rhythm. Methods that do not rely on input modifications to incorporate music theory are interesting but likely challenging due to changes required to the core Transformer implementation as was the case with the work of [HVU+18b] on relative positional encoding.
@@
Another key challenge is in bridging the gap between approaches more naturally formulated at the symbolic token level and approaches suitable to be used with acoustic tokens. As in other areas of structured audio processing any attempt to split audio into the underlying units (e.g. phones in speech or chords in music) is a highly non-trivial endeavour requiring expert annotators to act under the uncertainty of determining boundaries among those units. In the case of music, this is further compounded by the need to decide on the level of some of those units (e.g. degree of loudness). Thus, approaches capable of automatically inferring segmentations or marginalising over all possible segmentations would need to be developed.

### v2

@@ Abstract
In recent years, foundation models (FMs) such as large language models (LLMs) and latent diffusion models (LDMs) have profoundly impacted diverse sectors, including music. This comprehensive review examines state-of-the-art (SOTA) pre-trained models and foundation models in music, spanning from representation learning, generative learning and multimodal learning. We first contextualise the significance of music in various industries and trace the evolution of AI in music. By delineating the modalities targeted by foundation models, we discover many of the music representations are underexplored in FM development. Then, emphasis is placed on the lack of versatility of previous methods on diverse music applications, along with the potential of FMs in music understanding, generation and medical application. By comprehensively exploring the details of the model pre-training paradigm, architectural choices, tokenisation, finetuning methodologies and controllability, we emphasise the important topics that should have been well explored, like instruction tuning and in-context learning, scaling law and emergent ability, as well as long-sequence modelling, etc. A dedicated section presents insights into music agents, accompanied by a thorough analysis of datasets and evaluations essential for pre-training and downstream tasks. Finally, by underscoring the vital importance of ethical considerations, we advocate that following research on FM for music should focus more on such issues as interpretability, transparency, human responsibility, and copyright issues. The paper offers insights into future challenges and trends on FMs for music, aiming to shape the trajectory of human-AI collaboration in the music realm.
@@ III-A1 Traditional Music Information Retrieval Tasks
Given the extensive task list above, there is no single unified resource for providing task definitions and benchmark methodologies, although there exist textbooks that cover a subset of the above tasks (e.g. [Mül15, Ler22]), with ISMIR conference proceedings [^10] https://ismir.net/conferences/ being useful for providing additional technical details. Review and overview papers do exist for specific tasks (which might or might not be up to date with related literature depending on the publication year), including on onset detection [BDA+05], music transcription [BDDE19], chord estimation [MSRNDB14], melody extraction [SGER14], score following [OLS03], music source separation [CFL+18], music classification [FLTZ10], audio identification [CBKH05], cover song detection [SGH10], optical music recognition [CZJP20], plus general roadmaps for the community [SMB+13].
@@ III-A2 Multimodal Music Understanding Tasks
Music instruction following Mu-llama [LHSS23a] and MusiLingo [DML+24] integrated MERT[LYZ+24], a pre-trained music encoder, and Llama, a pre-trained natural language decoder. Mu-llama employs a Llama-adapter on the upper transformer layers for modality alignment [ZHL+23a] while MusiLingo utilises a simple linear projection on the initial transformer layer, yielding enhanced performance. Given the diverse range of tasks within the Music Understanding category, recent works have explored the application of Multimodal Instruction Tuning to enhance the performance of Music Language Models. LLark [GDSB23b] and JMLA [DYL+24] investigate how this approach can be effectively applied to the training of Music LLMs, aiming to improve their versatility and effectiveness across a wide spectrum of music-related tasks. Both models employ a similar architectural approach, utilising cross-attention modules to connect audio encoders with pre-trained LLMs. This integration allows for the effective processing of both audio and textual information. To create robust instruction-tuning datasets, these studies leverage traditional music tagging datasets, employing data augmentation techniques to expand them into comprehensive datasets suitable for fine-tuning Multimodal LLMs. By incorporating this approach, these models demonstrate the potential for more adaptive and comprehensive music understanding capabilities, effectively bridging the gap between audio processing and natural language understanding in the music domain.
@@ III-A3 Vocal Music Understanding
There have been several attempts to approach singer ID through representation learning. [LN19] proposed a triplet network that learns a joint embedding for monophonic and mixture singing. [ZQY+21] transformed the classification problem into a query searching task by replacing the softmax layer with a k-nearest neighbours (KNN) layer. [TLR23] trained SSL features on source-separated singing voices and compared them with speech SSL features in terms of classification accuracy.
@@ IV-F Foundation Models as Music Agents
Broadly speaking, an agent is a kind of entity that has desires, beliefs, intentions, and the ability to take actions [ZNAP95]. In the realm of AI, an agent is an artificial entity that can set objectives autonomously or semi-autonomously, perceive the environment, work out plans, reason and make decisions, and take actions to achieve these objectives, powered by the technologies of artificial intelligence [XCG+23]. In terms of music agents, they refer to AI agents dedicated to music and related domains, which can understand, generate, and convert music in symbolic, acoustic, and other modalities.
The development of AI agents has undergone a long process. Early attempts at building AI agents focus on improving the specific capabilities of an agent, such as sensing the environment, symbolic reasoning, taking actions with reinforcement learning, or handling specific tasks. However, the keys to AI agents are the general capabilities of reasoning, planning, perception, and action. While focusing on specific aspects of AI agents can result in steady progress, it might not be able to make fundamental enough improvements to push AI agents to practical usage.
Large language models (LLMs) that emerged in recent years have revolutionised the era of language and multimodality and brought strong capabilities in reasoning, planning, perception, and generation, making them a good fit as the core module of AI agents. Some pioneer investigations such as AutoGPT [Teab], HuggingGPT [SST+23a], and Visual ChatGPT [WYQ+23] leverage LLMs to understand user requests, reason and plan according to objectives, decompose complicated tasks into subtasks, revoke external tools for task execution, and response generation. On the one hand, LLM-based agents extend the capabilities of LLMs to handle complicated and/or multimodality tasks. On the other hand, LLMs empowered AI agents to a new level with strong general capabilities that have never been possible using traditional AI agent technologies. Due to the great potential of LLM-powered AI agents, a lot of research efforts have been made in this area, such as improving the reasoning and planning capabilities [WWS+22, Teaa], benchmarking AI agents [SST+23b], enhancing tool use [SDYD+23, YSC+24], designing agents to handle challenging tasks [SST+23a], and applying AI agents to more domains, including audio [LZL+23, LZL+24] and music [YSL+23]. In the next paragraph, we will review the typical AI agent in music and audio domains.
@@
  • Loop Copilot [ZMX+23] introduces an innovative system combining large language models with specialised AI music models to streamline the collaborative creation of music loops. This system utilises a conversational interface for dynamic, iterative music editing, and a global attribute table to ensure consistency throughout the creation process. Besides, it is not limited to creating music based on vague text inputs but allows for fine-grained musical edits, including adding or removing tracks and making localised adjustments to modes and tempos. This capability enhances the system’s utility in detailed music production tasks.
@@
Although there are some preliminary investigations on music agents, there are still a lot of space to improve. We introduce some future work of music agents bellow.
  • Improve reasoning and planning ability. The key capability of the music agents is enabled by the foundation model behind it. When handling complex tasks, the reasoning and planning ability of the foundation model plays a critical role in understanding the user request, decomposing the whole task into subtasks, organizing complicated working flows, and calling suitable models and tools for task execution [HC23, HGM+23]. Thus, improving the reasoning and planning capability has always been the top priority in building music agents as well as general artificial intelligence.
  • From semi-autonomous to full-autonomous agents. Nowadays, most music agents are semi-autonomous, requiring a human in the loop, such as initiating user requests, setting system configurations, providing suitable model and tool sets, and strong interactions in the agent workflows. While, Autonomous agents based on LLM are popular research directionss[HGM+23, WMF+24, YLL+24, YLY+24]. In order to play more important and critical roles in helping musicians and music consumers, music agents should evolve from semi-autonomous to full-autonomous, gradually reducing the degree of user interference when fulfilling complex user requests.
@@ IV-G Scaling Laws for Music
Scaling laws [KMH+20, HKK+20, HBM+22, AAA+23] characterise how the performance of large language models (LLMs) (usually measured with a cross-entropy loss on a held-out set) scales as a power law of model size, dataset size, and training compute. Scaling laws can predict the loss of LLMs with a given number of model parameters and data size (the number of training tokens, i.e., batch size $\times$ training steps) and can determine the optimal configuration of model size and number of training tokens given a computational budget. There are different formulations for scaling laws, such as OpenAI’s scaling laws [KMH+20], Chinchilla scaling laws [HBM+22], and data-constraint scaling laws [MRB+24]. OpenAI’s scaling laws are more precise, while Chinchilla’s laws are simplified based on some assumptions. For simplicity [^14] Strictly speaking, Chinchilla scaling laws have obvious errors. For example, given two models with model size $N_{1}$ and $N_{2}$ where $N_{1}>N_{2}$, when increasing the same amount of training tokens $\Delta D$, the loss reductions of the two models are the same according to Chinchilla laws as shown in Equation 2. However, it is evidently not the case in practice since the larger model $N_{1}$ will have more loss reduction than $N_{2}$. But for ease of explanation, we just use Chinchilla scaling laws as an example here., we briefly describe Chinchilla scaling laws as follows:
@@
There are several typical uses of scaling laws: 1) predict the loss of LLMs with a given model size and training tokens; 2) allocate the optimal model/data size given a compute budget; 3) determine the relative scaling of model and data with an increased compute budget; 4) determine the ratio of data to model given a compute budget.
Traditional scaling laws are mostly studied in language language models for text [KMH+20, HKK+20, HBM+22, AAA+23, MRB+24], which are not applicable for music. MuPT [QBM+24] is the first work to study the scaling laws for music, with dedicated design on symbolic music data representations and scaling laws in the data-constraint and over-fitting scenarios. While the great benefits of MuPT in guiding the modelling scaling of symbolic music generation, there is still a lack of scaling laws for music understanding and generation in the audio domain. Later, some studies emerged investigating scaling laws for multimodality [HKK+20, AYC+23] and especially for speech-language model [CM24]. Since multimodal data such as images, videos, and speech are high-dimensional and complex and not usually modelled by language models, it is not straightforward to study the scaling laws on multimodal data. Therefore, they usually tokenise multimodal data into tokens and model them as the next token prediction. In this way, they could study the scaling laws similarly to those in text-based large language models.
@@
Since there is not much research on scaling laws for music in the audio domain (although the scaling laws from multimodal and especially speech domain could provide much insight into the audio music domain, they are not directly applicable), and the emergent ability of music FMs is under-explored, we call for actions to proactively study the scaling properties of different data representations and model paradigms for audio music understanding and generation in the future. Such research topics include, but are not limited to, which type of downstream capability of large models can be inferred by the small models, how the properties of audio tokenisers relate to the parameters of scaling law, and what the minimal number of model parameters or training data is for a specific type of downstream tasks etc.
@@ IV-H Additional Future Improvements
In this subsection, we will shortly discuss potential research topics related to foundation models for music. Other than the foundation model techniques discussed in previous subsections, the following paragraphs will include domain knowledge, long sequence modelling and causal modelling. For discussions on fairness, transparency, and potential issues from model bias, please refer to subsectionVI-A2.
@@ IV-H1 Domain Knowledge
The model architecture described in this section so far, apart from small modifications, can be construed as ‘standard’ in the sense that it has been explored for a plethora of other tasks both somewhat linked with music (e.g. speech) as well as entirely different from it (e.g. machine translation). Although such an approach is quite appealing [HBL12], the limited data circumstances and inadequacies of architecture/optimisation in the music domain have led to a substantial interest in incorporating domain knowledge. This reversal towards an appreciation of domain expertise has brought about a revival in domain-specific engineering of machine learning approaches in many areas, including music [Ser13]. There are quite a few disciplines from which domain knowledge for music foundation models can be sourced. Acoustics, for example, gave rise to spectrum-based representations, which continue to find their use in deep learning architectures [DJP+20a]. The same applies to psychoacoustic chroma tokens [M1̈5] which can assist deep learning should it fail to automatically infer the importance of timbre information from the data [LYZ+24]. Music theory provides a further rich source of information which has inspired numerous approaches and will possibly continue to do so.
@@
One key challenge here is how to integrate music theory and related concepts into existing standard architectures rather than designing custom architectures to fit one or more music theories, as was done before [Wid98] and still remains popular in the area [HOB23], or uncover new ones by means of automatic theoretical analysis [YV16]. One popular direction for integrating music theory into the Transformer and related architectures is input modification. Although a simple approach, it provides opportunities for encoding a wide range of music-theoretical information. So far, the work in this direction has focused on encoding lower-level relative attributes, such as pitch interval, duration and onset [GKH23], and some higher-level attributes, such as bars [WERA16]. There remains to be an extension of this work to other higher-level relative attributes. The same applies to absolute attributes such as melody, harmony and rhythm. Methods that do not rely on input modifications to incorporate music theory are interesting but likely challenging due to changes required to the core Transformer implementation as was the case with the work of [HVU+18b] on relative positional encoding.
@@
Another key challenge lies in bridging the gap between approaches more naturally formulated at the symbolic token level and approaches suitable to be used with acoustic tokens. As in other areas of structured audio processing, any attempt to split audio into the underlying units (e.g. phones in speech or chords in music) is a highly non-trivial endeavour requiring expert annotators to act under the uncertainty of determining boundaries among those units. In the case of music, this is further compounded by the need to decide on the level of some of those units (e.g. degree of loudness). Thus, approaches capable of automatically inferring segmentations or marginalising over all possible segmentations would need to be developed.